OpenStudy (anonymous):

again this answer is incorrect i set the limits as -10

OpenStudy (amistre64):

 i assume itll look something like this than

OpenStudy (amistre64):

x=a..b y=0..f(x) z=-10..f(x,y)

OpenStudy (anonymous):

umm whats f(x) in this case?

OpenStudy (amistre64):

hmmm, y=x+4 in the xy plane

OpenStudy (amistre64):

and since the plane is horizontal to begin with; it transfers down to the z=-10 plane

OpenStudy (amistre64):

where does the y=o and y=x+4 planes intersect the 3rd plane?

OpenStudy (anonymous):

not quite shire

OpenStudy (amistre64):

urg!! my touchpad on the laptop is too sensitive, and goes flying off to wherever the cursor is hovering if i get to close to it :/

OpenStudy (amistre64):

x=-4..7 y=0..(x+4) z=4-2x-y

OpenStudy (amistre64):

\[\int_{-4}^7\int_{0}^{x+4}4-2x-y~dy~dx\]

OpenStudy (amistre64):

i forgot the -10 :) which is just a +10 on the function part

OpenStudy (amistre64):

\[\int_{-4}^7\int_{0}^{x+4}4-2x-y+10~dy~dx\] \[\int_{-4}^714(x+4)-2x(x+4)-\frac12(x+4)^2~dx\]

OpenStudy (anonymous):

im really confused on how came to this conclusion  isnt this the graph on the x y axis??

OpenStudy (amistre64):

in the xy plane yes, but the base of it all is still 10 below that

OpenStudy (anonymous):

i get the z limits but the limits for y and x have me confused could you explain em plx

OpenStudy (amistre64):

for the x, i just lowered it down to the base plane; y=0 and z=-10 2x+y+z = 4 2x+0-10=4 2x = 14; x=7 and of course it still has the y=x+4 limit in the z=-10 plane; y=0=x+4; x=-4

OpenStudy (amistre64):

but i think i see that if i keep it as such, i might need a split in the y

OpenStudy (amistre64):

2x+x+4 - 10 = 4 3x = 10 x=3' 1/3 seems to split there in the z=-10 plane; might have to think of doing the limits in another fashion eh

OpenStudy (anonymous):

hey there are many ways to do this i got the answer using my way incase you are curious=)

OpenStudy (amistre64):

as long as its the right answer, id like something to compare with yes :)

OpenStudy (amistre64):

ok, so my idea is, the y values change from 10/3 to 7 giving us a separate integration: \[\int_{-4}^{10/3}\int_{0}^{x+4}\int_{-10}^{4-2x-y}dzdydx+\int_{-4}^{10/3}\int_{0}^{-2x+14}\int_{-10}^{4-2x-y}dzdydx\] \[\int_{-4}^{10/3}\int_{0}^{x+4}14-2x-y~dydx+\int_{-4}^{10/3}\int_{0}^{-2x+14}14-2x-y~dydx\] \[\int_{-4}^{10/3}4x+54-2.5x^2~dx+\int_{-4}^{10/3} -42x+98+3x^2~dx\] and as slow as this computer is, i got no idea if thats right yet ..... better done on paper i spose

OpenStudy (amistre64):

yep, 295.777777